5. The total area of a regular tetrahedron is 110.85 m2, determine its base edge in m.
1
Expert's answer
2020-12-10T14:15:19-0500

Solution. A regular tetrahedron consists of four regular triangles. Therefore, the total area of a regular tetrahedron is equal to


"S_{total}=4\times\frac{1}{2}a^2sin60^0=a^2\sqrt{3}"

where a is the edge of the regular tetrahedron. Аccording to the condition of the problem


"S_{total}=110.85m^2"

Hense


"a^2\sqrt{3}=110.85""a=\sqrt{\frac{110.85}{\sqrt{3}}}\approx 8m"

Answer. 8m